all Systems Company, a telephone service and supply company, has just completed its fourth year of...

  1. all Systems Company, a telephone service and supply company, has just completed its fourth year of operations. The direct write-off method of recording bad debt expense has been used during the entire period. Because of substantial increases in sales volume and the amount of uncollectible accounts, the company is considering changing to the allowance method. Information is requested as to the effect that an annual provision of ½% of sales would have had on the amount of bad debt expense reported for each of the past four years. It is also considered desirable to know what the balance of Allowance for Doubtful Accounts would have been at the end of each year. The following data have been obtained from the accounts:

    Year of Origin of Accounts Receivable Written Off as Uncollectible
    Year Sales Uncollectible Accounts Written Off 1st 2nd 3rd 4th
    1st $1,140,000 $1,050 $1,050
    2nd 1,600,000 2,700 1,250 $1,450
    3rd 2,780,000 12,100 3,500 2,800 $5,800
    4th 3,700,000 18,150 4,150 6,150 $7,850

    Required:

    1. Assemble the desired data. Enter a decrease in the amount of expense as a negative number and all other amounts as positive numbers.

    Call Systems Company
    Bad Debt Expense
    Year Expense Actually Reported Expense Based on Estimate Increase (Decrease) in Amount of Expense Balance of Allowance Account, End of Year
    1st $ $ $ $
    2nd
    3rd
    4th

    2. Experience during the first four years of operations indicated that the receivables were either collected within two years or had to be written off as uncollectible. Does the estimate of ½% of sales appear to be reasonably close to the actual experience with uncollectible accounts originating during the first two years
Answer #1

1)

Bad debt expense
Year Expense actually reported [A] Expense based on estimate [B]] Increase/(decrease)in amount of expense [B-A] Balance of allowance account,end of year
1 1050 1140000*.5%=5700 4650 0+5700-1050= 4650
2 2700 1600000*.5%=8000 5300 4650+8000-2700= 9950
3 12100 2780000*.5%=13900 1800 9950+13900-12100= 11750
4 18150 3700000*.5%=18500 350 11750+18500-18150= 12100

#Balance of allowance account at end of year :Beginning balance+ bad debt expense-amount written off

2)No,The estimate of 1/2% (or .5%) of sales does not appears to be reasonable close to actual experience with uncollectible accounts originating during the first two years.

Amount written off Estimated based on sales
1 1050 5700
2 2700 8000